OpenStudy (anonymous):

If f(x) = x – 3 and g(x) = 2x – 4, find (f + g)(x).

OpenStudy (anonymous):

(f+g)(x)=f(x)+g(x)=x-3+2x-4

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

@sauravshakya from there do I solve for x to find the answer?

OpenStudy (anonymous):

U dont need to solve for x, Silmplify that

OpenStudy (anonymous):

and thats the overall answer for the whole equation? @sauravshakya

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

@sauravshakya thank you <3

OpenStudy (anonymous):

WELCOME